OpenStudy (anonymous):

how do you evaluate 9P3 in probability? The 9 and 3 are small*

OpenStudy (anonymous):

\({}_9P_3\) is used to denote the number of permutations of 9 objects taken 3 at a time. You can compute it as follows: \[3!\binom93=\frac{9!}{(9-3)!}=\frac{9\times8\times\cdots\times2\times1}{6\times5\times\cdots\times2\times1}\]